# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v5

@@ Abstract
Data-driven methods such as reinforcement and imitation learning have achieved remarkable success in robot autonomy. However, their data-centric nature still hinders them from generalizing well to ever-changing environments. Moreover, collecting large datasets for robotic tasks is often impractical and expensive. To overcome these challenges, we introduce a new self-supervised neuro-symbolic (NeSy) computational framework, imperative learning (IL), for robot autonomy, leveraging the generalization abilities of symbolic reasoning. The framework of IL consists of three primary components: a neural module, a reasoning engine, and a memory system. We formulate IL as a special bilevel optimization (BLO), which enables reciprocal learning over the three modules. This overcomes the label-intensive obstacles associated with data-driven approaches and takes advantage of symbolic reasoning concerning logical reasoning, physical principles, geometric analysis, etc. We discuss several optimization techniques for IL and verify their effectiveness in five distinct robot autonomy tasks including path planning, rule induction, optimal control, visual odometry, and multi-robot routing. Through various experiments, we show that IL can significantly enhance robot autonomy capabilities and we anticipate that it will catalyze further research across diverse domains.
@@ 1 Introduction
With the rapid development of deep learning (LeCun et al., 2015), there has been growing interest in data-driven approaches such as reinforcement learning (Zhu and Zhang, 2021) and imitation learning (Hussein et al., 2017) for robot autonomy. However, despite these notable advancements, many data-driven autonomous systems are still predominantly constrained to their training environments, exhibiting limited generalization ability (Banino et al., 2018; Albrecht et al., 2022).
@@
As a comparison, humans are capable of internalizing their experiences as abstract concepts or symbolic knowledge (Borghi et al., 2017). For instance, we interpret the terms “road” and “path” as symbols or concepts for navigable areas, whether it’s a paved street in a city or a dirt trail through a forest (Hockley, 2011). Equipped with these concepts, humans can employ spatial reasoning to navigate new and complex environments (Strader et al., 2024). This adaptability to novel scenarios, rooted in our ability to abstract and symbolize, is a fundamental aspect of human intelligence that is still notably absent in existing data-driven autonomous robots (Garcez et al., 2022; Kautz, 2022).
@@
Though implicit reasoning abilities have garnered increased attention in the context of large language models (LLMs) (Lu et al., 2023; Shah et al., 2023a), robot autonomy systems still encounter significant challenges in achieving interpretable reasoning. This is particularly evident in areas such as geometric, physical, and logical reasoning (Liu et al., 2023). Overcoming these obstacles and integrating interpretable symbolic reasoning into data-driven models, a direction known as neuro-symbolic (NeSy) reasoning, could remarkably enhance robot autonomy (Garcez et al., 2022).
@@
While NeSy reasoning offers substantial potential, its specific application in the field of robotics is still in a nascent stage. A key reason for this is the emerging nature of the NeSy reasoning field itself, which has not yet reached a consensus on a rigorous definition (Kautz, 2022). On one side, a narrow definition views NeSy reasoning as an amalgamation of neural methods (data-driven) and symbolic methods, which utilize formal logic and symbols for knowledge representation and rule-based reasoning. Alternatively, the broader definition expands the scope of what constitutes a “symbol”. In this perspective, symbols are not only logical terms but also encompass any comprehensible, human-conceived concepts. This can include physical properties and semantic attributes, as seen in approaches involving concrete equations (Duruisseaux et al., 2023), logical programming (Delfosse et al., 2023), and programmable objectives (Yonetani et al., 2021a). This broad definition hence encapsulates reasoning related to physical principles, logical reasoning, geometrical analytics, etc. In this context, many literatures exemplified NeSy systems, e.g., model-based reinforcement learning (Moerland et al., 2023), physics-informed networks (Karniadakis et al., 2021), and learning-aided tasks such as control (O’Connell et al., 2022), task scheduling (Gondhi and Gupta, 2017), and geometry analytics (Heidari and Iosifidis, 2024).
In this article, we explore the broader definition of NeSy reasoning for robot autonomy and introduce a self-supervised NeSy learning framework, which will be referred to as imperative learning (IL). It is designed to overcome the well-known issues of existing learning frameworks for robot autonomy: (1) Generalization ability: Many data-driven systems, including reinforcement learning models (Banino et al., 2018; Albrecht et al., 2022), remain largely confined to their training environments, displaying limited generalization capabilities. One reason for this limitation is their inability to learn explicit commonsense rules, which hinders their effective transfer to new environments. This encourages us to delve into symbolic reasoning techniques. (2) Black-box nature: The underlying causal mechanisms of data-driven models are mostly unknown, while a problematic decision can have catastrophic side effects in robotic tasks (Dulac-Arnold et al., 2021). Consequently, these models frequently encounter difficulties when applied to real-world situations, which further encourages us to delve into NeSy systems. (3) Label intensiveness: Labeling data for robotic tasks like imitation learning (Zare et al., 2023) often incurs higher costs than tasks in computer vision due to the reliance on specialized equipment over basic human annotations (Ebadi et al., 2022). For example, annotating accurate ground truth for robot planning is extremely complicated due to the intricacies of robot dynamics. This underscores the critical need for efficient self-supervised learning methods. (4) Sub-optimality: Separately training neural and symbolic modules can result in sub-optimal integration due to compounded errors (Besold et al., 2021). This motivates the investigation of an end-to-end NeSy learning approach.
IL is formulated to address the above issues with a single design. It is inspired by an interesting observation: while data-driven models predominantly require labeled data for parameter optimization, symbolic reasoning models are often capable of functioning without labels. Yet, both types of models can be optimized using gradient descent-like iterative methods. Take, for instance, geometrical models such as bundle adjustment (BA), physical models like model predictive control (MPC), and discrete models such as the A search over a graph. They can all be optimized without providing labels, even though they are analogous to data-driven models when formulated as optimization problems. IL leverages this property of both approaches by enforcing each method to mutually correct the other, thereby creating a novel self-supervised learning paradigm. To optimize the entire framework, IL is formulated as a special bilevel optimization (BLO), solved by back-propagating the errors from the self-supervised symbolic models into the neural models. The term “imperative” is adopted to highlight and describe this passive self-supervised learning procedure.
@@
This article is inspired by and built on our previous works in different fields of robot autonomy including local planning (Yang et al., 2023), global planning (Chen et al., 2024), simultaneous localization and mapping (SLAM) (Fu et al., 2024), feature matching (Zhan et al., 2024), and multi-agent routing (Guo et al., 2024). These works introduced a prototype of IL in several different domains while failing to introduce a systematic methodology for NeSy learning in robot autonomy. This article fills this gap by formally defining IL, exploring various optimization challenges of IL in different robot autonomy tasks, and introducing new applications such as rule induction and optimal control. Additionally, we introduce the theoretical background for solving IL based on BLO, propose several practical solutions to IL by experimenting with the five distinct robot autonomy applications, and demonstrate the superiority of IL over state-of-the-art (SOTA) methods in their respective fields.
@@ 2.1 Bilevel Optimization
Methodologically, gradient-based BLO can be generally divided into approximate implicit differentiation (AID), iterative differentiation (ITD), and value-function-based approaches. Based on the explicit form of the gradient (or hypergradient) of the upper-level objective function via implicit differentiation, AID-based methods adopt a generalized iterative solver for the lower-level problem as well as an efficient estimate for Hessian-inverse-vector product of the hypergradient (Domke, 2012; Pedregosa, 2016; Liao et al., 2018; Arbel and Mairal, 2022a). ITD-based methods approximate the hypergradient by directly taking backpropagation over a flexible inner-loop optimization trajectory using forward or backward mode of autograd (Maclaurin et al., 2015; Franceschi et al., 2017; Finn et al., 2017; Shaban et al., 2019; Grazzi et al., 2020). Value-function-based approaches reformulated the lower-level problem as a value-function-based constraint and solved this constrained problem via various constrained optimization techniques such as mixed gradient aggregation, log-barrier regularization, primal-dual method, and dynamic barrier (Sabach and Shtern, 2017; Liu et al., 2020a; Li et al., 2020a; Sow et al., 2022; Liu et al., 2021b; Ye et al., 2022). Recently, large-scale stochastic BLO has been extensively studied both in theory and in practice. For example, Chen et al. (2021) and Ji et al. (2021) proposed a Neumann series-based hypergradient estimator; Yang et al. (2021), Huang and Huang (2021), Guo and Yang (2021), Yang et al. (2021), and Dagréou et al. (2022) incorporated the strategies of variance reduction and recursive momentum; and Sow et al. (2021) developed an evolutionary strategies (ES)-based method without computing Hessian or Jacobian.
Theoretically, the convergence of BLO has been analyzed extensively based on a key assumption that the lower-level problem is strongly convex (Franceschi et al., 2018; Shaban et al., 2019; Liu et al., 2021b; Ghadimi and Wang, 2018; Ji et al., 2021; Hong et al., 2020; Arbel and Mairal, 2022a; Dagréou et al., 2022; Ji et al., 2022a; Huang et al., 2022). Among them, Ji and Liang (2021) further provided lower complexity bounds for deterministic BLO with (strongly) convex upper-level functions. Guo and Yang (2021), Chen et al. (2021), Yang et al. (2021), and Khanduri et al. (2021) achieved a near-optimal sample complexity with second-order derivatives. Kwon et al. (2023); Yang et al. (2024) analyzed the convergence of first-order stochastic BLO algorithms. Recent works studied a more challenging setting where the lower-level problem is convex or satisfies Polyak-Lojasiewicz (PL) or Morse-Bott conditions (Liu et al., 2020a; Li et al., 2020a; Sow et al., 2022; Liu et al., 2021b; Ye et al., 2022; Arbel and Mairal, 2022b; Chen et al., 2023; Liu et al., 2021c). More results on BLO and its analysis can be found in the survey (Liu et al., 2021a; Chen et al., 2022).
@@
2.2 Learning Frameworks in Robotics
We summarize the major learning frameworks in robotics, including imitation learning, reinforcement learning, and meta-learning. The others will be briefly mentioned.
@@ 2.2 Learning Frameworks in Robotics
has garnered significant attention recently, particularly with its application to training deep neural networks (Bengio et al., 1991; Thrun and Pratt, 2012). Unlike conventional learning approaches, meta-learning leverages datasets and prior knowledge of a task ensemble to rapidly learn new tasks, often with minimal data, as seen in few-shot learning. Numerous meta-learning algorithms have been developed, encompassing metric-based (Koch et al., 2015; Snell et al., 2017; Chen et al., 2020b; Tang et al., 2020; Gharoun et al., 2023), model-based (Munkhdalai and Yu, 2017; Vinyals et al., 2016; Liu et al., 2020b; Co-Reyes et al., 2021), and optimization-based methods (Finn et al., 2017; Nichol and Schulman, 2018; Simon et al., 2020; Singh et al., 2021; Bohdal et al., 2021; Zhang et al., 2024; Choe et al., 2024). Among them, optimization-based approaches are often simpler to implementation. They are achieving state-of-the-art results in a variety of domains.
@@ 2.3 Neuro-symbolic Learning
Physics-infused networks
(PINs) integrate physical laws directly into the architecture and training of neural networks (Raissi et al., 2019), aiming to enhance the model’s capability to solve complex scientific, engineering, and robotics problems (Karniadakis et al., 2021). PINs embed known physical principles, such as conservation laws and differential equations, into the network’s loss function (Duruisseaux et al., 2023), constraints, or structure, fostering greater interpretability, generalization, and efficiency (Lu et al., 2021). For example, in fluid dynamics, PINs can leverage the Navier-Stokes equations to guide predictions, ensuring adherence to fundamental flow properties even in data-scarce regions (Sun and Wang, 2020). Similarly, in structural mechanics, variational principles can be used to inform the model about stress and deformation relationships, leading to more accurate structural analysis (Rao et al., 2021).
In robot autonomy, PINs have been applied to various tasks including perception (Guan et al., 2024), planning (Romero et al., 2023), control (Han et al., 2024), etc. By embedding the kinematics and dynamics of robotic systems directly into the learning process, PINs enable robots to predict and respond to physical interactions more accurately, leading to safer and more efficient operations. Methodologies for this include but are not limited to embedding physical laws into network (Zhao et al., 2024), enforcing initial and boundary conditions into the training process (Rao et al., 2021), and designing physical constraint loss functions, e.g., minimizing an energy functional representing the physical system (Guan et al., 2024).
@@ 3.1 Structure
The proposed framework of imperative learning (IL) is illustrated in Figure 1, which consists of three modules, i.e., a neural system, a reasoning engine, and a memory module. Specifically, the neural system extracts high-level semantic attributes from raw sensor data such as images, LiDAR points, IMU measurements, and their combinations. These semantic attributes are subsequently sent to the reasoning engine, a symbolic process represented by physical principles, logical inference, analytical geometry, etc. The memory module stores the robot’s experiences and acquired knowledge, such as data, symbolic rules, and maps about the physical world for either a long-term or short-term period. Additionally, the reasoning engine performs a consistency check with the memory, which will update the memory or make necessary self-corrections. Intuitively, this design has the potential to combine the expressive feature extraction capabilities from the neural system, the interpretability and the generalization ability from the reasoning engine, and the memorability from the memory module into a single framework. We next explain the mathematical rationale to achieve this.

### v6

@@ Abstract
Data-driven methods such as reinforcement and imitation learning have achieved remarkable success in robot autonomy. However, their data-centric nature still hinders them from generalizing well to ever-changing environments. Moreover, labeling data for robotic tasks is often impractical and expensive. To overcome these challenges, we introduce a new self-supervised neuro-symbolic (NeSy) computational framework, imperative learning (IL), for robot autonomy, leveraging the generalization abilities of symbolic reasoning. The framework of IL consists of three primary components: a neural module, a reasoning engine, and a memory system. We formulate IL as a special bilevel optimization (BLO), which enables reciprocal learning over the three modules. This overcomes the label-intensive obstacles associated with data-driven approaches and takes advantage of symbolic reasoning concerning logical reasoning, physical principles, geometric analysis, etc. We discuss several optimization techniques for IL and verify their effectiveness in five distinct robot autonomy tasks including path planning, rule induction, optimal control, visual odometry, and multi-robot routing. Through various experiments, we show that IL can significantly enhance robot autonomy capabilities and we anticipate that it will catalyze further research across diverse domains.
@@ 1 Introduction
With the rapid development of deep learning (LeCun et al., 2015), there has been growing interest in data-driven approaches such as reinforcement learning (Zhu and Zhang, 2021) and imitation learning (Hussein et al., 2017) for robot autonomy. However, despite these notable advancements, many data-driven autonomous systems are still predominantly constrained to their training environments, exhibiting limited generalization ability (Banino et al., 2018; Albrecht et al., 2022). As a comparison, humans are capable of internalizing their experiences as abstract concepts or symbolic knowledge (Borghi et al., 2017). For instance, we interpret the terms “road” and “path” as symbols or concepts for navigable areas, whether it’s a paved street in a city or a dirt trail through a forest (Hockley, 2011). Equipped with these concepts, humans can employ spatial reasoning to navigate new and complex environments (Strader et al., 2024). This ability to apply spatial reasoning in unfamiliar contexts is a hallmark of human intelligence, one that remains largely missing from current data-driven autonomous robots (Garcez et al., 2022; Kautz, 2022).
@@
Though implicit reasoning abilities have garnered increased attention in the context of large language models (LLMs) (Lu et al., 2023; Shah et al., 2023a), robot autonomy systems still encounter significant challenges in achieving interpretable reasoning. This is particularly evident in tasks involving geometric, physical, and logical reasoning (Liu et al., 2023). Overcoming these obstacles and integrating interpretable symbolic reasoning into data-driven models, a direction known as neuro-symbolic (NeSy) reasoning, could remarkably enhance robot autonomy (Garcez et al., 2022).
@@
While NeSy reasoning offers substantial potential, its specific application in the field of robotics is still in a nascent stage. A key reason for this is the emerging nature of the NeSy reasoning field itself, which has not yet reached a consensus on a rigorous definition (Kautz, 2022). On one side, a narrow definition views NeSy reasoning as an amalgamation of neural methods (data-driven) and symbolic methods, which utilize formal logic and symbols for knowledge representation and rule-based reasoning. Alternatively, the broader definition expands the scope of what constitutes a “symbol”. In this perspective, symbols are not only logical terms but also encompass any comprehensible, human-conceived concepts. This can include physical properties and semantic attributes, as seen in approaches involving concrete equations (Duruisseaux et al., 2023), logical programming (Delfosse et al., 2023), and programmable objectives (Yonetani et al., 2021a). This broad definition hence supports reasoning in both continuous space (physical reasoning) and discrete space (logical reasoning). In this context, many literatures exemplified NeSy systems, e.g., model-based reinforcement learning (Moerland et al., 2023), physics-informed networks (Karniadakis et al., 2021), and learning-aided tasks such as control (O’Connell et al., 2022), task scheduling (Gondhi and Gupta, 2017), and geometry analytics (Heidari and Iosifidis, 2024).
In this article, we explore the broader definition of NeSy reasoning for robot autonomy and introduce a self-supervised NeSy learning framework, which will be referred to as imperative learning (IL). It is designed to be a unified framework following three principles: (1) Data-driven and Scalable: Self-Supervised. Large data-driven models have demonstrated promising performance in robotics. However, labeling data for robotic tasks is often significantly more costly than in computer vision, as it typically requires specialized equipment rather than simple human annotations (Ebadi et al., 2022). For instance, generating accurate ground truth for robot planning is particularly challenging due to the complexity of robot dynamics. As a result, to ensure the scalability of data-driven approaches in robot autonomy, the development of efficient self-supervised learning frameworks is essential. (2) Generalizable and Interpretable: Neuro-Symbolic Learning with Physical and Logical Structure. Robotic tasks often involve underlying structured knowledge governed by both physical laws and logical relationships (Dulac-Arnold et al., 2021), such as kinematic and logical constraints, task preconditions, and safety rules. Rather than relying purely on opaque neural models, IL integrates neural learning with symbolic reasoning components that encode these domain-specific structures. This hybrid design enables the system to make decisions that are easier to interpret, while also improving generalization to novel tasks and environments by leveraging prior knowledge about how the world works. (3) Modular and Global Optimal: End-to-end Trainable. A modular design reduces complexity by decomposing a system into independent yet interconnected components, making development, debugging, and interpretation more manageable. However, training neural and symbolic modules separately can lead to suboptimal integration, as errors may propagate and accumulate across components (Besold et al., 2021). To overcome this limitation, IL adopts a modular yet end-to-end trainable architecture, retaining the interpretability and flexibility of modular design, while enabling joint optimization across the entire pipeline.
IL is designed to embody the three core principles within a unified framework, inspired by a key observation: both neural models and symbolic methods can be formulated as optimization problems. The key distinction lies in their supervision requirements: neural models typically rely on labeled data for training, whereas symbolic reasoning models can often operate without explicit supervision. For instance, symbolic methods such as bundle adjustment (BA) for geometric reasoning, model predictive control (MPC) for physical reasoning, and A search for logical reasoning can all be framed as optimization problems and solved without labeled data. IL leverages this property by enabling neural and symbolic components to mutually refine one another. To achieve this, IL is formulated as a special bilevel optimization (BLO), where gradients are back-propagated from self-supervised symbolic modules into the neural networks. This design creates a novel self-supervised learning paradigm, in which symbolic reasoning imperatively guides the learning process. The term “imperative” is thus used to emphasize this passive form of self-supervised, neuro-symbolic, and end-to-end learning.
@@
This article is inspired by and built on our previous works in different fields of robot autonomy including local planning (Yang et al., 2023), global planning (Chen et al., 2024), simultaneous localization and mapping (SLAM) (Fu et al., 2024), feature matching (Zhan et al., 2024), and multi-agent routing (Guo et al., 2024). These works introduced a prototype of IL in several different domains while failing to introduce a systematic framework for NeSy learning in robot autonomy. This article fills this gap by formally defining IL, exploring various optimization challenges of IL in different robot autonomy tasks, and introducing new applications such as rule induction and optimal control. Additionally, we introduce the theoretical background for solving IL based on BLO, propose several practical solutions to IL by experimenting with the five distinct robot autonomy applications, and demonstrate the superiority of IL over state-of-the-art (SOTA) methods in their respective fields.
@@ 2.1 Bilevel Optimization
Methodologically, gradient-based BLO can be generally divided into approximate implicit differentiation (AID), iterative differentiation (ITD), and value-function-based approaches. Based on the explicit form of the gradient (or hypergradient) of the upper-level objective function via implicit differentiation, AID-based methods adopt a generalized iterative solver for the lower-level problem as well as an efficient estimate for Hessian-inverse-vector product of the hypergradient (Domke, 2012; Pedregosa, 2016; Liao et al., 2018; Arbel and Mairal, 2022a). ITD-based methods approximate the hypergradient by directly taking backpropagation over a flexible inner-loop optimization trajectory using forward or backward mode of autograd (Maclaurin et al., 2015; Franceschi et al., 2017; Finn et al., 2017; Shaban et al., 2019; Grazzi et al., 2020). Value-function-based approaches reformulated the lower-level problem as a value-function-based constraint and solved this constrained problem via various constrained optimization techniques such as mixed gradient aggregation, log-barrier regularization, primal-dual method, and dynamic barrier (Sabach and Shtern, 2017; Liu et al., 2020b; Li et al., 2020a; Sow et al., 2022; Liu et al., 2021b; Ye et al., 2022). Recently, large-scale stochastic BLO has been extensively studied both in theory and in practice. For example, Chen et al. (2021) and Ji et al. (2021) proposed a Neumann series-based hypergradient estimator; Yang et al. (2021), Huang and Huang (2021), Guo and Yang (2021), Yang et al. (2021), and Dagréou et al. (2022) incorporated the strategies of variance reduction and recursive momentum; and Sow et al. (2021) developed an evolutionary strategies (ES)-based method without computing Hessian or Jacobian.
Theoretically, the convergence of BLO has been analyzed extensively based on a key assumption that the lower-level problem is strongly convex (Franceschi et al., 2018; Shaban et al., 2019; Liu et al., 2021b; Ghadimi and Wang, 2018; Ji et al., 2021; Hong et al., 2020; Arbel and Mairal, 2022a; Dagréou et al., 2022; Ji et al., 2022a; Huang et al., 2022). Among them, Ji and Liang (2021) further provided lower complexity bounds for deterministic BLO with (strongly) convex upper-level functions. Guo and Yang (2021), Chen et al. (2021), Yang et al. (2021), and Khanduri et al. (2021) achieved a near-optimal sample complexity with second-order derivatives. Kwon et al. (2023); Yang et al. (2024) analyzed the convergence of first-order stochastic BLO algorithms. Recent works studied a more challenging setting where the lower-level problem is convex or satisfies Polyak-Lojasiewicz (PL) or Morse-Bott conditions (Liu et al., 2020b; Li et al., 2020a; Sow et al., 2022; Liu et al., 2021b; Ye et al., 2022; Arbel and Mairal, 2022b; Chen et al., 2023; Liu et al., 2021c). More results on BLO and its analysis can be found in the survey (Liu et al., 2021a; Chen et al., 2022).
@@
2.2 Robot Learning Frameworks
We summarize the major robot learning frameworks, including imitation learning, reinforcement learning, and meta-learning. The others will be briefly mentioned.
@@ 2.2 Robot Learning Frameworks
has garnered significant attention recently, particularly with its application to training deep neural networks (Bengio et al., 1991; Thrun and Pratt, 2012). Unlike conventional learning approaches, meta-learning leverages datasets and prior knowledge of a task ensemble to rapidly learn new tasks, often with minimal data, as seen in few-shot learning. Numerous meta-learning algorithms have been developed, encompassing metric-based (Koch et al., 2015; Snell et al., 2017; Chen et al., 2020b; Tang et al., 2020; Gharoun et al., 2023), model-based (Munkhdalai and Yu, 2017; Vinyals et al., 2016; Liu et al., 2020c; Co-Reyes et al., 2021), and optimization-based methods (Finn et al., 2017; Nichol and Schulman, 2018; Simon et al., 2020; Singh et al., 2021; Bohdal et al., 2021; Zhang et al., 2024; Choe et al., 2024). Among them, optimization-based approaches are often simpler to implementation. They are achieving state-of-the-art results in a variety of domains.
@@ 2.3 Neuro-symbolic Learning
Physical reasoning
in NeSy learning refers to approaches that integrate physical knowledge into the learning process to improve interpretability, generalization, and safety. Two major paradigms in this space are physics-informed and physics-infused learning. Physics-informed learning typically incorporates physical laws into the training objective by adding physics-based loss terms or constraints, guiding neural networks to produce physically consistent outputs (Raissi et al., 2019; Karniadakis et al., 2021). For example, in fluid dynamics, the Navier–Stokes equations can be enforced as soft penalties to regularize flow predictions (Sun and Wang, 2020), and in structural mechanics, variational formulations are used to encode stress-strain behavior (Rao et al., 2021). In contrast, physics-infused learning embeds physical laws or physics-based models directly into the model architecture (Vargas Venegas and Huang, 2023; Martin-Linares et al., 2023). For instance, Romero et al. (2023); Han et al. (2024) incorporated differentiable trajectory optimization into learning-based planning and control pipelines, allowing end-to-end learning with explicit physical reasoning; Zhao et al. (2024) programmed energy conservation laws into the model architecture for motion prediction in off-road driving. These methods tightly couple data-driven components with physical structure, enabling more expressive and interpretable models that can propagate gradients through physical interactions. We aim to incorporate physical reasoning via differentiable optimization within a unified neuro-symbolic learning framework for robot autonomy.
@@ 3.1 Structure
The proposed framework of imperative learning (IL) is illustrated in Figure 1, which consists of three modules, i.e., a neural learning system, a symbolic reasoning engine, and a memory module. Specifically, the neural system extracts high-level semantic attributes from raw sensor data such as images, LiDAR points, IMU measurements, and their combinations. These semantic attributes are subsequently sent to the reasoning engine, a symbolic process represented by physical principles, logical inference, analytical geometry, etc. The memory module stores the robot’s experiences and acquired knowledge, such as data, symbolic rules, and maps about the physical world for either a long-term or short-term period. Additionally, the reasoning engine performs a consistency check with the memory, which will update the memory or make necessary self-corrections. Intuitively, this design has the potential to combine the expressive feature extraction capabilities from the neural system, the interpretability and the generalization ability from the reasoning engine, and the memorability from the memory module into a single framework. We next explain the mathematical rationale to achieve this.